I AM BUYING A 70" DIAGANOL TV, WHAT IS THE LENGTH AND THE HIEGTH?
Answer by stanbon(75887)

I AM BUYING A 70" DIAGANOL TV, WHAT IS THE LENGTH AND THE HIEGTH?
The Aspect Ratio of wide-screen TV's is 16:9
---
Solve:
The width and length of your TV are 16x and 9x
---
Use Pytagoras to solve for "x":
(16x)^2 + (9x)^2 = 70^2
---
337x^2 = 4900
x^2 = 14.54
x = 3.8131
---
Width = 16x = 61"
Height = 9x = 31.3"
=========================
Cheers,
Stan H.
=========================